OpenStudy (anonymous):

Find the measure of an arc given that the shaded region has an area of 45.55 square centimeters if the radius is 6?

OpenStudy (anonymous):

What is area of arc?

OpenStudy (anonymous):

that shows the diagram

OpenStudy (anonymous):

arc is the same as the angle given

OpenStudy (anonymous):

there is no angle

OpenStudy (anonymous):

Actually, lets start with the area of the circle

OpenStudy (anonymous):

whats 45.55

OpenStudy (anonymous):

the area of shaded region

OpenStudy (anonymous):

do you find area of circle and put 45.55 over area as a proportion to x/360? to find arc angle?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Then you have x/360 and times this by circumference to get arc length

OpenStudy (anonymous):

is it 144.99?

OpenStudy (anonymous):

Area of circle: 36pi sector area/circle area: 45.55/36pi ~ 0.403 Circumference: 12pi Arc length: 12pi*45.55/36pi = 45.55/3 ~15.18

OpenStudy (anonymous):

area of full circle is : \[\pi r ^{2}\] the proportion of a sector to the full area is same as the angle of the sector to the full angle: \[45.55/\pi*6 ^{2}=x/2 \pi\] from here : \[x=(45.55*2)/36\]

OpenStudy (anonymous):

notice that x is in radians. To get the answer in centimiters multiply it by 6